OpenStudy (anonymous):

For the pair of functions, f and g, find (f o g)(x) and (g o f)(x) f(x)=4-7x g(x)=5x+5

OpenStudy (anonymous):

(f o g)(x)=-31-35x (g o f)(x)=25-35x

OpenStudy (sidsiddhartha):

hint: \[FoG(X) = f(g(x))\] and \[GoF(x) = g(f(x))\]

OpenStudy (sidsiddhartha):

i'm doing the first one: \[(f o g)(x)=f[5x+5]=4-7(5x+5)\] like this got that?